# Finding separatrices of dynamical flows with Deep Koopman Eigenfunctions

### v1

@@ Abstract
Many natural systems, including neural circuits involved in decision making, can be modeled as high-dimensional dynamical systems with multiple stable states. While existing analytical tools primarily describe behavior near stable equilibria, characterizing separatrices – the manifolds that delineate boundaries between different basins of attraction – remains challenging, particularly in high-dimensional settings. Here, we introduce a numerical framework leveraging Koopman Theory combined with Deep Neural Networks to effectively characterize separatrices. Specifically, we approximate Koopman Eigenfunctions (KEFs) associated with real positive eigenvalues, which vanish precisely at the separatrices. Utilizing these scalar KEFs, optimization methods efficiently locate separatrices even in complex systems. We demonstrate our approach on synthetic benchmarks, ecological network models, and recurrent neural networks trained on neuroscience-inspired tasks. Moreover, we illustrate the practical utility of our method by designing optimal perturbations that can shift systems across separatrices, enabling predictions relevant to optogenetic stimulation experiments in neuroscience.
@@ 1 Introduction
Recurrent neural networks (RNNs) are widely used both in neuroscience, as models of circuit dynamics, and in machine learning, as powerful tools for sequential data processing [1, 2]. A key goal in neuroscience is to reverse-engineer these models in order to understand the underlying dynamical mechanisms [1, 2]. In particular, many cognitive tasks such as decision-making [3] and associative memory [4] can be modeled as multistable dynamical systems, where distinct decisions or memories correspond to different stable attractor states in phase space. Transitions between these attractors are governed by the geometry of the basins of attraction and, crucially, by the separatrix: the manifold that delineates the boundary between basins (Figure 1A).
A reverse-engineering method that has yielded significant insights about RNN computations involves finding approximate fixed points and linearising around them [5]. This involves minimizing a scalar function – the kinetic energy $q(x)=\|f(x)\|^{2}$ – to locate these points (Figure 1B). Once found, the linearisation of the dynamics at the fixed point can shed light on the mechanism of computations [6, 7, 8, 9, 10, 11, 12].
@@
Ideally, we would have a scalar function analogous to the kinetic energy – smooth, yet vanishing precisely on the separatrix (Figure 1C). This would allow gradient-based optimization to locate the decision boundary, help visualize the geometry of this manifold, and enable the design of optimal decision-changing perturbations (Figure 1A).
In this work, we propose a novel method to characterize separatrices in high-dimensional black-box dynamical systems by leveraging Koopman operator theory [13]. Specifically, we approximate scalar-valued Koopman eigenfunctions (KEFs) with positive real eigenvalues using deep neural networks. These eigenfunctions vanish precisely on the separatrix. Because they are scalar functions, gradient-based optimization can be used to efficiently trace out the separatrix, even in high dimensions.
We apply this framework to synthetic systems, ecological models, and RNNs trained on neuroscience-inspired tasks. In addition, we demonstrate that the learned KEFs can be used to design minimal perturbations that push the system across separatrices—a setting relevant to experimental protocols such as optogenetic stimulation.
@@
  • We empirically demonstrate the method on systems of increasing complexity: from 1D synthetic models to 64-dimensional trained RNNs, and an 11D ecological model.
@@ 2 Related Work
Koopman theory has recently been used to evaluate similarity between dynamical systems, both in neuroscience [14]—where it is applied to study the temporal structure of computation—and in machine learning, where it has been used to compare training dynamics across models [15]. These approaches typically analyze system-level behavior using dynamic mode decomposition [16, 17, 18] a finite-dimensional approximation of the Koopman Operator.
In parallel, deep learning methods have emerged as powerful tools for solving partial differential equations (PDEs). Notably, the Deep Ritz Method [19] and Deep Galerkin Method (DGM) [20] demonstrate how deep neural networks (DNNs) can approximate solutions to variational and differential problems. A related line of work uses physics-informed neural networks (PINNs), which incorporate known PDE constraints as part of the loss function during training [21].
Koopman-based embeddings have also been proposed as a tool for analyzing the internal dynamics of RNNs. In [22], the authors show that eigenvectors of finite-dimensional approximations of the Koopman operator can uncover task-relevant latent structure in RNNs. More generally, several works explore DNN-based approximations of Koopman operators for learning meaningful embeddings of nonlinear dynamics [23, 24, 25].
Finally, our approach is conceptually connected to work on the geometry of Koopman eigenfunctions themselves. In particular, [26] studies the level sets of KEFs and their relationship to isostables and isochrons in systems with stable fixed points. These constructions motivate our interest in identifying scalar functions that vanish on separatrices, as a means of understanding boundaries between basins of attraction.
@@ 3.1 KEFs as Scalar Separatrix Indicators
where the $\dot{\square}$ is shorthand for the time derivative $\frac{d}{dt}\square$ and $f:\mathcal{X}\to\mathcal{X}$ defines the dynamics on an $N$ dimensional state space $\mathcal{X}$.
Our goal is to construct a smooth scalar function $\psi$ that vanishes precisely on the separatrix between basins of attraction. A natural candidate for such a function is the inverse of the time it takes for a trajectory to reach an attractor. Near the separatrix, this time diverges, and hence the inverse goes to zero. Mathematically, we want an observable $\psi(\boldsymbol{x}(t))$ of the state of the system $\boldsymbol{x}$ that keeps growing as time evolves and $\boldsymbol{x}$ converges to an attractor of the system. This motivates a function $\psi(\boldsymbol{x})$ that, when $\boldsymbol{x}(t)$ is a trajectory of the dynamics, satisfies:
$\psi(\boldsymbol{x}(t))=\psi(\boldsymbol{x}(0))e^{\lambda t}.$ (2)
This is precisely the behavior of a Koopman eigenfunction (KEF) with eigenvalue $\lambda>0$. Note that Koopman eigenfunctions are usually introduced in a different manner, and the supplementary material shows the connection to our description.
A Koopman eigenfunction $\psi:\mathcal{X}\to\mathbb{R}$ satisfies:
$\frac{d}{dt}\bigg{(}\psi\big{(}\boldsymbol{x}(t)\big{)}\bigg{)}=\lambda\psi(\boldsymbol{x}(t))\quad\Leftrightarrow\quad\nabla\psi(\boldsymbol{x})\cdot f(\boldsymbol{x})=\lambda\psi(\boldsymbol{x}).$ (3)
This is the Koopman Partial differential equation (PDE) (see Appendix C for its relationship to the Koopman Operator). $\lambda$ relates to the timescale of $\psi$ and is an important hyperparameter of our method (Appendix I).
@@ A Koopman eigenfunction $\psi:\mathcal{X}\to\mathbb{R}$ satisfies:
where $p(\boldsymbol{x})$ is a sampling distribution over the phase space [19, 20]. As with any eigenvalue problem, this loss admits the trivial solution $\psi\equiv 0$. To discourage such solutions, we introduce a shuffle-normalization loss where the two terms are sampled independently from the same distribution:
@@
Figure 2: Our method to approximate KEFs in three bistable systems. (A) A 1D system $\dot{x}=x-x^{3}$. The curve shows $f(x)$, and its fixed points in light green – ‘o’s stable and ‘x’s unstable. (B) The kinetic energy $q(x)$ of this system. (C) the true KEF (8) and it’s DNN approximation obtained by our method. (D,E,F) Damped Duffing Oscillator in 2D (G,H,I) 2-unit GRU [27] RNN trained on 1-bit flop flop (1BFF) [5] and our KEFs.
@@
The system has three fixed points, corresponding to minima of $q(x)$ (Figure 2B). A $\lambda=1$ KEF can be derived analytically (Appendix A)):
@@ 3.2 Challenges and Solutions
We observe this behavior empirically in Appendix Figure 8, where independently initialized networks converge to different spurious solutions. To mitigate this, we introduce a balance regularization term that biases $\psi$ to have nonzero values in opposing basins, encouraging sign changes across the separatrix. Specifically, we define:
@@
where $A(x)$ and $B(y)$ are KEFs to the respective 1D problems. Crucially, when $\mu=1$, the eigenfunction depends only on $x$ and ignores $y$—capturing only the separatrix along $x=0$.
Figure 3 (bottom) illustrates this effect: different values of $\mu$ yield KEFs aligned with different separatrices. To address this, we train multiple KEFs $\{\psi_{i}(x)\}_{i=1}^{k}$ and combine them via the geometric mean:
$\psi_{\text{combined}}(x)=\left(\prod_{i=1}^{k}\psi_{i}(x)\right)^{1/k},$ (13)
which preserves the zero-level sets of all $\psi_{i}$.
To encourage each $\psi_{i}$ to capture different separatrix components, we propose sampling from multiple distributions centered somewhere on the separatrix. First, we obtain points on the separatrix by interpolating between fixed points and identifying the transition via binary search and numerical integration of the ODE (1). Around each such point $\beta_{j}$, we define a local distribution $\mathcal{N}(\beta_{j},\sigma_{j}^{2}I)$, using a range of scales $\{\sigma_{j}\}$ to span both fine and global structure. For each distribution, we compute losses $\mathcal{L}^{j}_{\text{ratio}}$ and $\mathcal{L}^{j}_{\text{bal}}$, and minimize the weighted sum:
This procedure biases each KEF to resolve local separatrix geometry while retaining global consistency.
@@ 3.3 Demonstrations
We first demonstrate our method on a low-dimensional recurrent neural network trained to perform a two-bit flip flop (2BFF) task. Specifically, we use a 3-unit gated recurrent unit (GRU) network [27]. The trained network exhibits four stable fixed points (Figure 4), corresponding to different memory states of the task.
To overcome the degeneracies described in Figure 3, we adopt a targeted sampling strategy. We first identify points on the separatrix by interpolating between pairs of fixed points and performing binary search: at each step, we simulate the dynamics to determine basin membership and refine the search. Around these discovered separatrix points, we construct concentric isotropic Gaussian distributions, and sample from them to train on the loss $\mathcal{L}_{\text{total}}$, (14).
@@
We next apply our method to a high-dimensional ecological model: a generalized Lotka–Volterra (gLV) system fit to genus-level abundance data from a mouse model of antibiotic-induced Clostridioides difficile infection (CDI) [28]. The system has five stable fixed points. Following [29] we focus our analysis to two of these fixed points representing healthy and diseased microbial states.
We optimize the KEF in the full 11-dimensional state space. For interpretability, we follow the projection approach of [29], visualizing the dynamics in the 2D plane spanned by the two chosen stable fixed points and the origin (see Figure 5). Although the KEF is trained entirely in the original 11-dimensional space, its zero level set (light green curve) aligns well with the true separatrix (orange line) computed using a grid of initial conditions in the 2D slice [29].
Figure 5: KEF approximation in a fitted 11D gLV model of CDI [28, 29]. Zero level set of the KEF aligns with the separatrix in a 2D projection plane.
@@
We test our method in a setting where there are no fixed points along the separatrix. We construct a system which oscillates at a fixed frequency ($\dot{\theta}=1$), but converges to one of two preferred amplitudes ($\dot{r}=(r-2)-(r-2)^{3}$). The system has three limit-cycles, two of them stable ($r=1,3$) and one unstable ($r=2$). In Figure 6B we visualise the flow, it’s kinetic energy and the limit cycles. The system has no fixed points, and thus fixed point analysis is futile. We utilize Radial basis function neural network [30] to parameterise the KEF (Appendix F).
@@
64D RNN Solving 1BFF
To test our method in a high-dimensional setting, we train a 64-unit vanilla RNN to solve a one-bit flip flop (1BFF) task. The resulting dynamics exhibit bistability between two memory states. As before, we identify a point on the separatrix via binary search along the line connecting the fixed point attractors, simulating the dynamics to classify basin membership at each step.
We then train the KEF network using samples drawn from isotropic Gaussian distributions centered at this separatrix point. Due to the high dimensionality, direct visualization of the KEF and the dynamics is not feasible. Instead, we devise a curve-based validation approach.
We construct multiple Hermite polynomial curves that interpolate between the two stable fixed points. The curvature of each curve is parameterized by a random vector, and each is defined by a parameter $\alpha\in[0,1]$, where $\alpha=0$ corresponds to one attractor and $\alpha=1$ to the other (see appendix E). Because each curve continuously connects the fixed points, it must cross the separatrix. Figure 7A shows a 2D PCA projection of several such Hermite curves. Crucially, the actual curves span the entire 64D space. We simulate dynamics from 100 points along each curve and determine their final basin to infer where each curve crosses the separatrix, forming a ground-truth reference.
@@
Finally, we demonstrate how the KEF can be used to design minimal perturbations that shift the state across the separatrix (similar to Figure 1A, Figure 4C). Given a base point $x_{\text{base}}$, we aim to solve:
$x^{*}=\operatorname*{arg\,min}_{x}\|x-x_{\text{base}}\|_{2}^{2}\quad\text{subject to}\quad|\psi(x)|=0,$ (15)
using the Adam optimizer [31] to minimize the relaxed loss:
$L(x)=|\psi(x)|+\gamma\|x-x_{\text{base}}\|_{2}^{2},$ (16)
with random initialization around $x_{\text{base}}$. Figure 7D compares the distance from $x_{\text{base}}$ to our optimal point versus separatrix points identified along Hermite curves and the direct line connecting the base point to the destination fixed point. Our method yields the shortest valid perturbation.
Figure 7: Validation of KEF approximation in a 64D RNN trained on 1BFF. (A) PCA projection of Hermite curves between fixed points coloured by true basin labels. (B) KEF values along three Hermite curves versus curve parameter $\alpha$, as well the true separatrix point along the curve. (C) Comparison between true and predicted separatrix positions along curves. (D) Distance from $x_{\text{base}}$ to perturbation targets. The KEF-guided solution yields the smallest perturbation crossing the separatrix.

### v2

@@ Abstract
Many natural systems, including neural circuits involved in decision making, are modeled as high-dimensional dynamical systems with multiple stable states. While existing analytical tools primarily describe behavior near stable equilibria, characterizing separatrices—the manifolds that delineate boundaries between different basins of attraction—remains challenging, particularly in high-dimensional settings. Here, we introduce a numerical framework leveraging Koopman Theory combined with Deep Neural Networks to effectively characterize separatrices. Specifically, we approximate Koopman Eigenfunctions (KEFs) associated with real positive eigenvalues, which vanish precisely at the separatrices. Utilizing these scalar KEFs, optimization methods efficiently locate separatrices even in complex systems. We demonstrate our approach on synthetic benchmarks, ecological network models, and high-dimensional recurrent neural networks trained on either neuroscience-inspired tasks or fit to real neural data. Moreover, we illustrate the practical utility of our method by designing optimal perturbations that can shift systems across separatrices, enabling predictions relevant to optogenetic stimulation experiments in neuroscience. Code available on GitHub.
@@ 1 Introduction
Recurrent neural networks (RNNs) are widely used both in neuroscience, as models of circuit dynamics, and in machine learning, as powerful tools for sequential data processing [barak_recurrent_2017, vyas_computation_2020]. A key goal in neuroscience is to reverse-engineer these models in order to understand the underlying dynamical mechanisms [barak_recurrent_2017, vyas_computation_2020]. In particular, many cognitive tasks such as decision-making [machens_flexible_2005] and associative memory [hopfield_neural_1982] can be modeled as multistable dynamical systems, where distinct decisions or memories correspond to different stable attractor states in phase space. Transitions between these attractors are governed by the geometry of the basins of attraction and, crucially, by the separatrix: the manifold that delineates the boundary between basins (Figure 1A).
A reverse-engineering method that has yielded significant insights about RNN computations involves finding approximate fixed points and linearising around them [sussillo_opening_2013]. This involves minimizing a scalar function—the kinetic energy $q(x)=\|f(x)\|^{2}$—to locate these points (Figure 1B). Once found, the linearisation of the dynamics at the fixed point can shed light on the mechanism of computations [carnevale_dynamic_2015, maheswaranathan_reverse_2019, maheswaranathan_universality_2019, finkelstein_attractor_2021, mante_context-dependent_2013, liu_encoding_2024, driscoll_flexible_2024, low_remapping_2023, chaisangmongkon_computing_2017, pagan_individual_2025, wang_flexible_2018].
@@
Ideally, we would have a scalar function analogous to the kinetic energy—smooth, yet vanishing precisely on the separatrix (Figure 1C). This would allow gradient-based optimization to locate the decision boundary, help visualize the geometry of this manifold, and enable the design of optimal decision-changing perturbations (Figure 1A).
In this work, we propose a novel method to characterize separatrices in high-dimensional black-box dynamical systems by leveraging Koopman operator theory [koopman_hamiltonian_1931, budisic_applied_2012]. Specifically, we approximate scalar-valued Koopman eigenfunctions (KEFs) with positive real eigenvalues using deep neural networks. These eigenfunctions vanish precisely on the separatrix. Because they are scalar functions, gradient-based optimization can be used to efficiently trace out the separatrix, even in high dimensions.
We apply this framework to synthetic systems, ecological models, RNNs trained on neuroscience-inspired tasks, and trained to reproduce neural recordings. In addition, we demonstrate that the learned KEFs can be used to design minimal perturbations that push the system across separatrices—a setting relevant to experimental protocols such as optogenetic stimulation.
@@
  • We empirically demonstrate the method on systems of increasing complexity: from low-dimensional synthetic models to 668-dimensional RNNs fit to mouse neural data.
@@ 2 Related Work
Koopman theory has recently been used to evaluate similarity between dynamical systems, both in neuroscience [ostrow_beyond_2023]—where it is applied to study the temporal structure of computation—and in machine learning, where it has been used to compare training dynamics across models [redman_identifying_2024]. These approaches typically analyze system-level behavior using dynamic mode decomposition [schmid_dynamic_2010, tu_dynamic_2014, brunton_koopman_2016], a finite-dimensional approximation of the Koopman operator.
In parallel, deep learning methods have emerged as powerful tools for solving partial differential equations (PDEs). Notably, the Deep Ritz Method [e_deep_2018] and Deep Galerkin Method (DGM) [sirignano_dgm_2018] demonstrate how deep neural networks (DNNs) can approximate solutions to variational and differential problems. A related line of work uses physics-informed neural networks (PINNs), which incorporate known PDE constraints as part of the loss function during training [raissi_physics-informed_2019].
Koopman-based embeddings have also been proposed as a tool for analyzing the internal dynamics of RNNs. In [naiman_operator_2023], the authors show that eigenvectors of finite-dimensional approximations of the Koopman operator can uncover task-relevant latent structure in RNNs. More generally, several works explore DNN-based approximations of Koopman operators for learning meaningful embeddings of nonlinear dynamics [lusch_deep_2018, yeung_learning_2019, deka_koopman-based_2022].
An alternative line of research for identifying Lagrangian Coherent Structures (LCS) employs the Finite-Time Lyapunov Exponent (FTLE) [haller_lagrangian_2000, tanaka_separatrices_2009, haller_lagrangian_2015], which quantifies sensitivity to initial conditions by measuring the exponential rate of separation between nearby trajectories over a finite time horizon. Ridges in the FTLE field reveal stable and unstable LCS, with the latter corresponding to separatrices in our terminology. These methods are most often applied to two- or three-dimensional fluid flows, where they delineate dynamically distinct regions in the velocity field.
Finally, our approach is conceptually connected to work on the geometry of Koopman eigenfunctions themselves. In particular, [mauroy_isostables_2013] studies the level sets of KEFs and their relationship to isostables and isochrons in systems with stable fixed points. In the setting of linear systems, and nonlinear systems topologically conjugate to them, [mezic_applications_2015, mezic_koopman_2021] establish links between KEF level sets and separatrices (stable manifolds). Together, these studies motivate our approach of deep-learning KEFs, as a method for identifying separatrices in general high-dimensional, multi-stable systems.
@@ 3.1 KEFs as Scalar Separatrix Indicators
Figure 2: Mapping the high dimensional dynamics of $\boldsymbol{x}$ with an unstable manifold – the separatrix – to the one dimensional linear dynamics of $\psi$ with instability at $\psi=0$, i.e., $\lambda>0$. We approximate the mapping $\psi(\boldsymbol{x})$ with a DNN.
@@
[…]
Our goal is to construct a smooth scalar function $\psi$ that vanishes precisely on the separatrix between basins of attraction. Consider two such basins (Figure 2). All we care about is the existence of the separatrix manifold, with dynamics moving away from it. The simplest such dynamics is a one-dimensional linear system $\dot{\psi}=\lambda\psi$, with $\lambda>0$. This motivates a mapping that projects $\boldsymbol{x}\in\mathcal{X}$ to $\psi(\boldsymbol{x})\in\mathbb{R}$, and induces these dynamics. Such a mapping needs to satisfy:
$\frac{d}{dt}\bigg(\psi\big(\boldsymbol{x}(t)\big)\bigg)=\lambda\psi\big(\boldsymbol{x}(t)\big)$ (2)
along any trajectory $\boldsymbol{x}(t)$ in $\mathcal{X}$. See Appendix J for a formal link between $\psi$ and the separatrix.
This is precisely the behavior of a Koopman eigenfunction (KEF) with eigenvalue $\lambda>0$. Note that Koopman eigenfunctions are usually introduced in a different manner, and Appendix C shows the connection to our description.
Equation (2) can be re-written as:
$\nabla\psi(\boldsymbol{x})\cdot f(\boldsymbol{x})=\lambda\psi(\boldsymbol{x}).$ (3)
by employing the chain-rule of differentiation and requiring it to hold at all $\boldsymbol{x}\in\mathcal{X}$. This is the Koopman Partial differential equation (PDE). $\lambda$ relates to the timescale of $\psi$ and is an important hyperparameter of our method (Appendix I).
@@
where $p(\boldsymbol{x})$ is a sampling distribution over the phase space [e_deep_2018, sirignano_dgm_2018]. As with any eigenvalue problem, this loss admits the trivial solution $\psi\equiv 0$. To discourage such solutions, we introduce a shuffle-normalization loss where the two terms are sampled independently from the same distribution:
@@
Figure 3: Our method to approximate KEFs in three bistable systems. (A) A 1D system $\dot{x}=x-x^{3}$. The curve shows $f(x)$, and its fixed points in light green – ‘o’s stable and ‘x’s unstable. (B) The kinetic energy $q(x)$ of this system. (C) the true KEF (8) and its DNN approximation obtained by our method. (D,E,F) Damped Duffing Oscillator in 2D (G,H,I) 2-unit GRU [chung_empirical_2014] RNN trained on 1-bit flip flop (1BFF) [sussillo_opening_2013] and our KEFs.
@@
The system has three fixed points, corresponding to minima of $q(x)$ (Figure 3B). A $\lambda=1$ KEF can be derived analytically (Appendix A):
@@ 3.2 Challenges and Solutions
We observe this behavior empirically in Appendix D, where independently initialized networks converge to different spurious solutions. To mitigate this, we introduce a balance regularization term that biases $\psi$ to have nonzero values in opposing basins, encouraging sign changes across the separatrix. Specifically, we define:
@@
where $A(x)$ and $B(y)$ are KEFs to the respective 1D problems. For example, when $\mu=1$, the eigenfunction depends only on $x$ and ignores $y$ – therefore unable to capture $y$-dependent separatrices. Figure 4 (bottom) illustrates this effect: different values of $\mu$ yield KEFs aligned with different separatrices.
Even in non-separable systems, this degeneracy can arise. Optimizing $\mathcal{L}_{\text{total}}$ can lead to a KEF that identifies some separatrices and ignores others (Appendix D). To address this, we train multiple KEFs $\{\psi_{i}(x)\}_{i=1}^{k}$, while using their input distributions to bias each one to capture a different separatrix. For $\psi_{i}(x)$, we choose two points in different basins of attraction, and then use a binary search on the line connecting them to find a point on the separatrix. Note that the KEF is still needed to obtain the full separatrix, and not just a point. Around each such point $\beta_{i}$, we define a local distribution $\mathcal{N}(\beta_{i},\sigma_{ij}^{2}I)$, using a range of scales $\{\sigma_{ij}\}_{j=1}^{J}$ to span both fine and global structure. For each distribution, we minimize the sum $\sum_{j=1}^{J}\mathcal{L}^{j}_{\text{total}}$. We then consider the union of the separatrices obtained from each of the KEFs to complete the picture (see 2-bit flip flop demonstration below).
@@ 3.3 Demonstrations
We first demonstrate our method on a low-dimensional recurrent neural network trained to perform a two-bit flip flop (2BFF) task. Specifically, we use a 3-unit gated recurrent unit (GRU) network [chung_empirical_2014]. The trained network exhibits four stable fixed points (Figure 5), corresponding to different memory states of the task.
To overcome the degeneracies described in Figure 4, we adopt a targeted sampling strategy. We first identify points on the separatrix by interpolating between pairs of fixed points and performing binary search: at each step, we simulate the dynamics to determine basin membership and refine the search. Around these discovered separatrix points, we construct concentric isotropic Gaussian distributions, and sample from them to train on the loss $\mathcal{L}_{\text{total}}$ (Appendix G).
@@
Figure 6: KEF approximation in a fitted 11D gLV model of CDI [stein_ecological_2013, jones_steady-state_2019]. Zero level set of the KEF aligns with the separatrix in a 2D projection plane.
@@
We next apply our method (Appendix G) to a high-dimensional ecological model: a generalized Lotka–Volterra (gLV) system fit to genus-level abundance data from a mouse model of antibiotic-induced Clostridioides difficile infection (CDI) [stein_ecological_2013]. The system has five stable fixed points. Following [jones_steady-state_2019] we focus our analysis to two of these fixed points representing healthy and diseased microbial states.
We optimize the KEF in the full 11-dimensional state space. For interpretability, we follow the projection approach of [jones_steady-state_2019], visualizing the dynamics in the 2D plane spanned by the two chosen stable fixed points and the origin (see Figure 6). Although the KEF is trained entirely in the original 11-dimensional space, its zero level set (light green curve) aligns well with the true separatrix (orange line) computed using a grid of initial conditions in the 2D slice [jones_steady-state_2019].
@@
We test our method in a setting where there are no fixed points along the separatrix. We construct a system which oscillates at a fixed frequency ($\dot{\theta}=1$), but converges to one of two preferred amplitudes ($\dot{r}=(r-2)-(r-2)^{3}$). The system has three limit-cycles, two of them stable ($r=1,3$) and one unstable ($r=2$). In Figure 7B we visualise the flow, its kinetic energy and the limit cycles. The system has no fixed points, and thus fixed point analysis is futile. We utilize Radial basis function neural network [powell_radial_1987] to parameterise the KEF (Appendix F).
@@
668D RNN fit to mouse neural activity
To demonstrate our method in a high-dimensional (see Appendix H for scaling results) and neuroscientifically relevant setting, we applied it to a recurrent neural network (RNN) trained to reproduce mouse neural activity from finkelstein_attractor_2021. The trained RNN exhibits bistability between two memory states. As in lower-dimensional systems, we first located a point on the separatrix by performing a binary search along the line connecting the fixed-point attractors, simulating the dynamics at each step to determine basin membership.
In the original experiment, mice were trained to respond to optogenetic stimulation of their sensory cortices, and the RNN was fit to the peristimulus time histogram of recorded neural activity. The network undergoes a bifurcation from monostability to bistability as a function of an external ramping input $u_{\text{ext}}$. For analysis, we fixed $u_{\text{ext}}=0.9$ within the bistable regime and trained the Koopman eigenfunction (KEF) network (Appendix G) using samples drawn from isotropic Gaussian distributions centered at the separatrix point.
Because of the high dimensionality, direct visualization of the learned KEF and dynamics is not feasible. Instead, we validated the model using a curve-based evaluation approach. We construct multiple Hermite polynomial curves that interpolate between the two stable fixed points. The curvature of each curve is parameterized by a random vector, and each is defined by a parameter $\alpha\in[0,1]$, where $\alpha=0$ corresponds to one attractor and $\alpha=1$ to the other (see appendix E). Because each curve continuously connects the fixed points, it must cross the separatrix. Figure 8A shows a 2D PCA projection of several such Hermite curves. Crucially, the actual curves span the entire 64D space. We simulate dynamics from 100 points along each curve and determine their final basin to infer where each curve crosses the separatrix, forming a ground-truth reference.
@@
Finally, we demonstrate how the KEF can be used to design minimal perturbations that shift the state across the separatrix (similar to Figure 1A, Figure 5C). In general, this involves an input-driven dynamics $\boldsymbol{\dot{x}}=\tilde{f}(\boldsymbol{x},\boldsymbol{u})$, with time-varying inputs $\boldsymbol{u}(t)$. To demonstrate the utility of the method we study a specific, simplified scenario in which the input is a strong instantaneous perturbation $\boldsymbol{u}(t):=\boldsymbol{\Delta}\delta(t)$, which moves the state $\boldsymbol{x}(0_{+})=\boldsymbol{x}(0_{-})+\boldsymbol{\Delta}$, after which the dynamics evolves according to $f(\boldsymbol{x}):=\tilde{f}(\boldsymbol{x},0)$.
Given a base point $\boldsymbol{x}(0_{-}):=\boldsymbol{x}_{\text{base}}$, just before perturbation, we aim to solve:
$\boldsymbol{\Delta}^{*}=\operatorname*{arg\,min}_{\Delta}\|\boldsymbol{\Delta}\|_{2}^{2}\quad\text{subject to}\quad|\psi(\boldsymbol{x_{\text{base}}}+\boldsymbol{\Delta})|=0.$ (13)
Taking advantage of the differentiable nature of $\psi$ in our method, we use the Adam optimizer [kingma_adam_2017] to find $\boldsymbol{\Delta}^{*}$ from random initialization. Figure 8D shows that indeed the optimised perturbation $\boldsymbol{\Delta}^{*}$ is smaller, compared to simply aiming the perturbation towards the target fixed point, or towards random points on the separatrix.
Figure 8: Validation of KEF approximation in a 668D RNN fit to mouse neural activity [finkelstein_attractor_2021]. (A) PCA projection of Hermite curves between fixed points coloured by true basin labels and separatrix points along the curves (red crosses). (B) KEF values along three Hermite curves versus curve parameter $\alpha$, as well the true separatrix point along the curve. (C) Comparison between true and predicted separatrix positions along curves. (D) Perturbation amplitudes $\|\boldsymbol{\Delta}\|$, i.e, distance from $x_{\text{base}}$ to perturbation targets. The KEF-guided solution yields the smallest perturbation crossing the separatrix.